Fermi National Laboratory
 

Operation Summary for April 5, 2010 - April 7, 2010

 

Monday April 5

The day shift began with Operations monitoring the Tevatron (TeV) store 7727 (86.49E30), the Recycler stash (251.0E10), and the Antiproton Source stack (17.0E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7727 at 3:04 PM. Cryo system technicians began system maintenance and TeV personnel began proton studies. An Accumulator RF station (ARF1) tripped off several time during the shift; Pbar personnel investigated and fixed the problem. The H- Source dome lost power for 15 minutes due to a controls issue.

During the evening shift, Operations put the TeV into HEP shot setup at 4:28 PM.

Operations established store 7729 at 5:27 PM with an average initial luminosity (AIL) of 353.6E30. Operations disabled the MTest beam at 6 PM. ARF1 tripped off several more times. At 8:08 PM, a lightning strike took down several devices; no beam was lost.

Tuesday April 6

The midnight shift began with Operations monitoring store 7729 (135.97E30) and the stash (190.3E10), with Pbar stacking (14.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:11 AM. ARF1 tripped off several more times. A couple of minor Booster RF trips occurred.

During the day shift, Operations terminated store 7729 at 9:38 AM and put the TeV into shot setup. Operations established store 7730 at 10:54 AM with an AIL of 313.8E30. At 10:18 AM, Operations halted beam for accesses into the Booster, Main Injector, and NuMI enclosures. The driving factor for the access was to fix a LCW leak in the Main Injector. Additionally, mechanical support personnel found an air leak on a beam valve in Recycler.

During the evening shift, the MI LCW work continued until 6:44 PM and the Recycler beam valve was repaired by 9:14 PM. Operations restored beam to the users by 10:22 PM. At 11:11 PM, MiniBooNE reported having controls problems that inhibited their beam; they consulted experts and decided to remain off through the night and begin repairs on Wednesday morning.

Wednesday April 7

The midnight shift began with Operations monitoring store 7730 (75.0E30) and the stash (54.0E10), with Pbar stacking (9.0E10), and with NuMI taking beam.

During the midnight shift, operators performed extensive Booster tuning all shift to improve beam positions. Operations established beam to MTest at 5:30 AM.

The Plans for Wednesday

The plans for today are stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/07/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory